OpenStudy (anonymous):

Solve the equation 4(x^2)y+8y=2x for y as a function of x.

OpenStudy (anonymous):

8y=2x-4x^2 y=x/4-x^2/2

OpenStudy (anonymous):

Factor y out like so: y[4(x^2) + 8] = 2x. Then, divide [4(x^2) +8] on both sides. Your answer will be: y= 2x/[4x^2 +8] which will simplify to y= 1/2x + 1/4x

OpenStudy (anonymous):

Djblank was really close, just forgot the y at the end of 4x^2

OpenStudy (anonymous):

I had y=1/2x+1/4x but it was wrong but then I left it as 2x/(4x^2+8) and it was right.

OpenStudy (anonymous):

Thanks Nick!